BBO Discussion Forums: The level you prefer - BBO Discussion Forums

Jump to content

  • 2 Pages +
  • 1
  • 2
  • You cannot start a new topic
  • You cannot reply to this topic

The level you prefer If you knew all cards

Poll: 6 or 7 ? (29 member(s) have cast votes)

6 or 7 ?

  1. 6NT (3 votes [10.34%])

    Percentage of vote: 10.34%

  2. 7NT (26 votes [89.66%])

    Percentage of vote: 89.66%

Vote Guests cannot vote

#21 User is offline   dcvetkov 

  • PipPipPipPip
  • Group: Full Members
  • Posts: 260
  • Joined: 2005-September-19
  • Location:Toronto, Canada
  • Interests:bridge, swimming, tennis

Posted 2008-April-22, 14:44

jdonn, on Apr 22 2008, 11:49 AM, said:

Why have cold sweats when the finesse is clearly percentage over the drop? You just finesse and pray :)

Is this really true? is it more likley then squeezing RHO? I know everyone is saying, yeah we make when queen is onside. But would you finesse at trick 11?
RHO can hold Qxx♥ and J♣ and is squeezed at trick 10, he has to come to doubleton heart.

I think you have to make some assumptions who has the last club (J♣), based on discards and table feel.
I agree its a pretty good 7, but trick 11,,,, grrr;)
[COLOR=blue] aka Dimitar
0

#22 User is offline   dcvetkov 

  • PipPipPipPip
  • Group: Full Members
  • Posts: 260
  • Joined: 2005-September-19
  • Location:Toronto, Canada
  • Interests:bridge, swimming, tennis

Posted 2008-April-22, 14:54

Opps

Edit

Transportation is a problem. Big. I would like to check if J♣ is dropping. cash the ace of hearts, run my winners AND end up in south hand, and then decide finesse or squeeze but I can not do that. Maybe too keen on squeeze endings.:)

So you need to take ♥ finesse early, before cashing your dummy winners. You may go down 2, but hey, tomorrow is another day.

Disregard my previous posts.
[COLOR=blue] aka Dimitar
0

#23 User is offline   OleBerg 

  • PipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 1,950
  • Joined: 2008-April-05
  • Gender:Male
  • Location:Copenhagen
  • Interests:Model-Railways.

Posted 2008-April-22, 15:26

dburn, on Apr 22 2008, 03:04 PM, said:

OleBerg, on Apr 22 2008, 02:51 PM, said:

gnasher, on Apr 22 2008, 01:35 PM, said:

Even if I have to play it single dummy, I think it's a good enough grand slam to bid at IMPs.  The two main chances are:

♣Jxx = 0.622 *3/7 = 0.267
♣Jx = 0.305 * 2/7 = 0.087
♣J singleton = 0.068 * 1/7 = 0.009
Subtotal = 0.363

♥Q onside = (1-0.363) / 2 = 0.318

So I've already got to 68%, without including either singleton ♥Q offside or the showup squeeze when LHO has ♣Jxxxx and ♥xxx.

You dont need all those calculations. Assuming they always bid 6nt at the other table, it's just good enough without the ♣10. Only Just.

Best Regards

Ole Berg

I'm not sure I understand this. Without the ♣10, you will make 7NT when West holds ♥Q, when East holds it singleton, and when East is void in hearts, all of which comes to about 57%.

That would not be good enough odds if there were only one other table (at IMPs - BAM is a different matter). Experience suggests that it is not good enough odds at matchpoints either, unless playing in a small field of uniformly high standard. One could imagine a number of pairs agreeing hearts at an early stage in the auction, finding out that ♥Q is missing, and settling for a small slam in hearts. Without ♣10, I would be content with 6NT in most fields at matchpoints (the form of scoring specified in the original post).

Even with ♣10, I still consider that many pairs will find it difficult to bid to 7NT, and I would estimate that 6NT making seven would score pretty well.

You're quite right, I miscalculated. I thuoght there were nine hearts in the combined hands. :P

Actually it only adds up to 55,6552%. Nowhere near the needed 58% (still assuming they always bid at least 6nt at the other table).

Can I have my ♣10 back, please?

Best Regards

Ole Berg
_____________________________________

Do not underestimate the power of the dark side. Or the ninth trumph.

Best Regards Ole Berg

_____________________________________

We should always assume 2/1 unless otherwise stated, because:

- If the original poster didn't bother to state his system, that means that he thinks it's obvious what he's playing. The only people who think this are 2/1 players.


Gnasher
0

#24 User is offline   han 

  • Under bidder
  • PipPipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 11,797
  • Joined: 2004-July-25
  • Gender:Male
  • Location:Amsterdam, the Netherlands

Posted 2008-April-22, 16:03

It's matchpoints, so anything more than 50% should do in a strong field.
Please note: I am interested in boring, bog standard, 2/1.

- hrothgar
0

#25 User is offline   gnasher 

  • Andy Bowles
  • PipPipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 11,993
  • Joined: 2007-May-03
  • Gender:Male
  • Location:London, UK

Posted 2008-April-22, 16:09

Where does 58% come from?

If you bid a grand slam and it makes, you gain 500, or 11 IMPs. If you bid a grand slam and go down one, you lose 1040, or 14 IMPs. Your break-even point is 14/(11 + 14), or 56%.
... that would still not be conclusive proof, before someone wants to explain that to me as well as if I was a 5 year-old. - gwnn
0

#26 User is offline   han 

  • Under bidder
  • PipPipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 11,797
  • Joined: 2004-July-25
  • Gender:Male
  • Location:Amsterdam, the Netherlands

Posted 2008-April-22, 16:12

Oh ok, I didn't read carefully enough.
Please note: I am interested in boring, bog standard, 2/1.

- hrothgar
0

#27 User is offline   gnasher 

  • Andy Bowles
  • PipPipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 11,993
  • Joined: 2007-May-03
  • Gender:Male
  • Location:London, UK

Posted 2008-April-22, 16:13

han, on Apr 22 2008, 11:03 PM, said:

It's matchpoints, so anything more than 50% should do in a strong field.

It was matchpoints, but we're talking about IMPs.
... that would still not be conclusive proof, before someone wants to explain that to me as well as if I was a 5 year-old. - gwnn
0

#28 User is offline   OleBerg 

  • PipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 1,950
  • Joined: 2008-April-05
  • Gender:Male
  • Location:Copenhagen
  • Interests:Model-Railways.

Posted 2008-April-23, 12:47

gnasher, on Apr 22 2008, 05:09 PM, said:

Where does 58% come from?

If you bid a grand slam and it makes, you gain 500, or 11 IMPs.  If you bid a grand slam and go down one, you lose 1040, or 14 IMPs.  Your break-even point is 14/(11 + 14), or 56%.

The 58% comes from my bad memory :)

I guess it is because I use it as a guideline, to make up for all the smaal quirks a contract often have. Things like, not being able to handle a 6-0 break.

Incidently, the suit-combination I thought I saw, a nine card fit with AKJT, is the "worst" allowable combination, you can have in a critical suit, when you bid a grand. (The suit is 57.91% likely to come in.)

To my knowledge at least, maybe Im wrong again. :)

Best Regard

Ole Berg
_____________________________________

Do not underestimate the power of the dark side. Or the ninth trumph.

Best Regards Ole Berg

_____________________________________

We should always assume 2/1 unless otherwise stated, because:

- If the original poster didn't bother to state his system, that means that he thinks it's obvious what he's playing. The only people who think this are 2/1 players.


Gnasher
0

  • 2 Pages +
  • 1
  • 2
  • You cannot start a new topic
  • You cannot reply to this topic

1 User(s) are reading this topic
0 members, 1 guests, 0 anonymous users